# Equivalence Checking of Quantum Circuits via Intermediary Matrix Product Operator

### v1

@@ VII.2 Scaling of Equivalent Circuits
VII.3 Scaling of Non-Equivalent Circuits
Here, we evaluate the scaling behavior for detecting non-equivalence between circuits. Since the ZX method cannot prove non-equivalence, we focus on comparing the MPO- and DD-based methods. Various errors of different intensity, as described in Sec. VII.1 and analyzed in Sec. VII.1, were injected into linear entanglement circuits. For the MPO-based method, the high SVD threshold $s_{\text{max}}=10^{-1}$ was used to detect non-equivalence. The results are presented in Fig. 7.
For all error types and severities, the DD method scales exponentially, likely because it becomes maximally large in all cases—an expected outcome given that it also scales exponentially for the equivalent case. In contrast, the MPO method scales polynomially, though more severe errors cause worse scaling, as expected due to the QMA-completeness of the equivalence-checking problem. These results can be extrapolated to understand the non-equivalence scaling when applied to SCA and full entanglement patterns, where we expect small errors to scale similarly to the equivalent case, and more severe to approach exponential scaling.
Interestingly, for circuits with low-severity errors, the MPO method’s scaling closely mirrors that of the linear equivalent case in Sec. VII.2. This shows that the MPO method surpasses previous methods by far for detecting small degrees of non-equivalence.

### v2

@@ VII.3 Scaling of equivalent circuits
VII.4 Scaling of non-equivalent circuits
Here, we evaluate the scaling behavior for detecting non-equivalence between circuits. Since the ZX method cannot prove non-equivalence, we focus on comparing the MPO- and DD-based methods. Various errors of different intensity, as described in Sec. VII.2 and analyzed in Sec. VII.2, were injected into linear entanglement circuits. For the MPO-based method, the high SVD threshold $s_{\text{max}}=10^{-1}$ was used to detect non-equivalence. The results are presented in Fig. 9.
For all error types and severities, the DD method scales exponentially, likely because it becomes maximally large in all cases—an expected outcome given that it also scales exponentially for the equivalent case. In contrast, the MPO method scales polynomially, though more severe errors cause worse scaling, as expected due to the QMA-completeness of the equivalence-checking problem. These results can be extrapolated to understand the non-equivalence scaling when applied to SCA and full entanglement patterns, where we expect small errors to scale similarly to the equivalent case, and more severe to approach exponential scaling. Concretely, by combining the results in this section with the error plots in Sec. VII.2, we can estimate that the transition from polynomial to exponential scaling of runtime as a function of system size occurs approximately for phase errors $\theta_{\text{error}}>10^{-2}\pi$ or when the number of swapped gates reaches $n_{\text{swap}}>5$. While additional gate errors do worsen scaling behavior, the tests performed here do not exhibit true exponential scaling. We expect such scaling to emerge only for significantly more non-equivalent circuits as the removed gates increases $|g|_{\text{removed}}\gg 10$.
Interestingly, for circuits with low-severity errors, the MPO method’s scaling closely mirrors that of the linear equivalent case in Sec. VII.3. This shows that the MPO method surpasses previous methods by far for detecting small degrees of non-equivalence.